OpenStudy (goformit100):

2013^2−1 equals (A) 2012×2014 (B) 2012×2013 (C) 2013×2014 (D) 2012×2012 (E) 2000×2013

OpenStudy (anonymous):

Expand the expression. \(2013^2-1=2013\times2013-1\)

OpenStudy (anonymous):

\[2013^2-1=2013^2-1^2=(2013+1)(2013-1)=(2014)(2012)\]

OpenStudy (goformit100):

Y I got the answer

OpenStudy (goformit100):

*ya